OpenStudy (anonymous):

Solve the equation 15(x - 3) = 12x - 43. Show your work.

OpenStudy (anonymous):

The answer is positive I know that

OpenStudy (anonymous):

15x-45=12x-43 15x-12x=-43+45 3x=2 x=3/2